40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 14974     Entry time: Thu Oct 17 11:19:28 2019     In reply to: 14973     Reply to this: 14983
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking activity last night 
  1. Tuning the MICH-->PRM output matrix element
    • Locked the PRMI with the carrier field resonant in the PRC.
    • REFL11 used to control PRCL, AS55 for MICH.
    • Turned on the sensing notches in the control filter bank. Drove a line in MICH at 311.10 Hz.
    • Tweaked the MICH-->PRM matrix element to minimize the coupling witnessed.
    • As shown in Attachment #1, the minimum coupling was found to be at the value -0.34 (the old value was -0.2655).
    • The minimum was very sharp. A 1% change from the optimum value increased the peak height by > x2. Is this reasonable?
  2. Some sensing matrix measurements: After tuning the output matrix element, I locked the PRMI (ETMs misaligned) in four configurations:
    • PRMI locked with carrier resonant. REFL11_I used for PRCL control, AS55_Q used for MICH control.
    • PRMI locked with sidebands resonant. REFL11_I used for PRCL control, AS55_Q used for MICH control.
    • PRMI locked with sidebands resonant. REFL11_I used for PRCL control, REFL165_I used for MICH control (based on sensing matrix measurement and offsets from previous config).
    • PRMI locked with sidebands resonant. REFL33_I used for PRCL control, AS55_Q used for MICH control.
    • The attached GIF shows the evolution of the demodulated sensing lines as we move through configurations.
       
    • The actual PDFs are attached as a zip, Attachment #2.
  3. PRMI locking with arms under ALS control
    • The arm cavity lengths were controlled as usual with ALS. This system needs some noise budgeting.
    • I set the CARM offset to -8 (arbitrarily chosen, approximately equal to 20nm, but anyways well above the cavity linewidth).
    • Then I re-aligned the PRM, and attemped to lock the PRMI using the 3f settings determined with no arm cavities --> no success.
    • Tried locking using the 1f error signals instead - in this config, the lock could be established.
    • However, I saw that there was significant light on the AS camera, and I had to put in an offset into the MICH loop to make ASDC go as low as possible.
    • I guess it is possible that the ALS control wasn't precise enough and the leaked light to the dark port was because of differential reflectivity of the arm cavities?
    • Anyways, I ran a sensing matrix measurement with the interferometer in this configuration, and I found that the MICH signal in REFL165 had rotated significantly.
    • I also found that the 3f DC offsets in this configuration were ~5x greater than what was the case for the lock with no arm cavities.

This is as far as I got last night. The first step is to see how reliable the settings determined last night are, today. I don't understand how changing the output matrix element can have brought about such a significant change in the MICH/PRCL separation in all the RF photodiodes.

Attachment 1: MICH2PRCLnulling.pdf  6 kB  Uploaded Thu Oct 17 12:24:46 2019  | Hide | Hide all
MICH2PRCLnulling.pdf
Attachment 2: consolidatedSensingMatrices.pdf.zip  672 kB  Uploaded Thu Oct 17 13:22:30 2019
ELOG V3.1.3-